OpenStudy (anonymous):

You draw two cards at random from a standard deck. What is the probability of drawing at least one diamond?

OpenStudy (anonymous):

A. 9/34 B. 3/4 C. 9/16 D. 15/34

OpenStudy (anonymous):

thank god I"ve been stuck on this question for 2 hours its very hard for me

OpenStudy (anonymous):

so 12/54 right?

hartnn (hartnn):

nopes how u got that?

hartnn (hartnn):

P(atleast 1 diamond) = 1- P(No diamond)

OpenStudy (anonymous):

someone joined earlier and sais that

hartnn (hartnn):

P (diamond) = 3/4

OpenStudy (anonymous):

ok

hartnn (hartnn):

u get how P(diamond) = 3/4?

OpenStudy (anonymous):

so u think c

OpenStudy (anonymous):

one guy said d

OpenStudy (anonymous):

but i do get it more then normally

OpenStudy (anonymous):

i dont tihnk its 3/4

hartnn (hartnn):

why?

OpenStudy (anonymous):

it was wrong

OpenStudy (anonymous):

but thxs for actually helping me

hartnn (hartnn):

i never said 3/4 is correct i just asked you why you think its not correct? :P

hartnn (hartnn):

P(No Diamond) = (3/4)*(38/51) = 19/34

hartnn (hartnn):

P(atleast 1 diamond) = 1-19/34 =...